2 csc2x + 4 = 1
2 csc2x + 3 = 0
??

Answer by stanbon(48558)

2 csc2x + 4 = 1
2 csc(2x) = -3
csc(2x) = -3/2
2x = arccsc(-3/2) = -41.81 degrees
x = -20.09 degrees
or
x = 339.09 degrees
============================
2 csc2x + 3 = 0
2csc(2x) = -3
csc(2x) = -3/2
Same answer as abvove
============================
Cheers,
Stan H.